OpenStudy (anonymous):

hi.. can anyone explain to me about how to find a parabola graph that matches to a certain given equation. Does anyone know what formula to use

OpenStudy (amistre64):

it easier if you have an equation to step thru

OpenStudy (anonymous):

These are parts of transformations: you already know that y=x^2 is the parabola y=(x-a)^2 + b is the transformations where b is the change in the y axis and a is the change in the x axis (note the negative)

OpenStudy (anonymous):

so if you had say y=(x-4)^2 + 9 then instead of the bottom of the parabola being (0,0) it would be shifted up the y axis 9 units and shifted right on the x axis 4 units

OpenStudy (anonymous):

i cant post the graph so i cant really post the problem because the question is the graph (to Amistre)

OpenStudy (mathteacher1729):

Sparky, have you tried graphing your problem? As in -- graphing y = (x-a)^2 + b for various values of a and b?

OpenStudy (anonymous):

no but ill try it

OpenStudy (mathteacher1729):

Geogebra can do this easily. Here is a quick sketch I made. :)

OpenStudy (anonymous):

count how many units it moves along the x axis and plug that into the equation as a then do the same thing for the y axis and plug that into the equation as b

OpenStudy (anonymous):

wowz thats alotta parabolas ~.~

OpenStudy (anonymous):

if you have the equation that you is thinking is right, enter it into wolframalpha.com and it will graph it for you and you can examine the two

OpenStudy (mathteacher1729):

Haha, yeah I had the "trace" feature on. If you don't trace, it won't do that. It'll just move the graph without leaving a trail. :)